Eating banana enable a person to perform about 40,000 J of work.  Assuming 25 % efficient of the human body to process the energy content of the banana, to what height does eating a banana enable a 60 kg person to climb?  


1
Expert's answer
2021-02-16T06:05:07-0500

"Work" "done""(w.d)=" "40,000J"

"Efficiency" "(E)" "=" "25"%

"mass""(m)" "=60kg"


"solution"


"w.d=f*d"

considering that the efficiency of human body is "25%"%,

hence therefore, the work ouput is given by:


"w.d=25\/100*40000"


"=10,000J"

Now from the formula of the work done


"w.d=f*d"

"however," "the" "person" "is" "climbing" "against" "the" "gravity"


where "force=mass" "*acceration" "due" "to" "gravity"


"hence"

"w.d=m*a*d"

"where" "d=height(h)"


"10,000=60*9.8*h"


"h=10,000\/588"


"h=17.01m"